AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 01:28:13 2025
Author
*
:
Subject
*
:
02.02 DAQ continues G22-86/R15_21 alpha background DSSSD leakage current = attachment 1 DSSSD #2 transient c. few hundred nA continue FEE64 temperatures OK - attachment 2 ADC data item stats - attachment 3 rates dominated by 'hot' channels aida01 ( 70k ) and aida15 ( 11k ) - all other FEE64s c. 0Hz per FEE64 Rate and Stat spectra - attachments 4-5 02.07 All stats zero'd 05.52 DAQ continues G22-86/R15_28 alpha background FEE64 temperatures OK - attachment 6 ADC data item stats - attachment 7 rates dominated by 'hot' channels aida01 ( 72k ), aida15 ( 11k ), aida12 ( 3k ) and aida05 ( 600 ) all other FEE64s c. 0Hz per FEE64 Stat spectra - attachment 8 aida12 asic #2 HEC channels high rate Merger and TapeServer - attachments 9-10 DSSSD leakage current = attachment 11 DSSSD #2 transient c. few hundred nA continue online per DSSSD 2D x-y spectra - attachnent 12 05.56 All histograms & stats zero'd 2x Check ASIC load ADC data item stats - attachment 7 rates dominated by 'hot' channels aida01 ( 72k ), aida15 ( 11k ) all other FEE64s c. 0Hz 10:28 DAQ continues G22-88/R15 FEE temperatures OK ADC data items stats dominated by aida01 (86k) and aida15 (9k) rest of stats approx OHz. Rates dominated by 2 hot channels per FEE Leakage current plot (attachment 15) System wide checks OK Merger OK (500,000 items per sec), Tape server OK (710 kb/sec being written) DSSD XY plot (attachment 16) - some noise coming through in certain strips on DSSD 3 DSSD Ey energy histograms (attachment 17) DSSD Ey vs Ex plots (attachment 18) - statistics for DSSD0,1,3 are consistent (DSSD 2 stats appear higher - unknown reason as yet) 11:10 Leakage current for DSSD1 (I2 on leakage current plot) remained 'turbulent'. We dropped the voltage from 150V to 100V. Leakage current dropped from ~4.5uA to ~4.2 uA. We will see if it remains more stable. Rates in DSSD1 increased during change in voltage so all histograms were zero'd and monitoring code reset. Transient also observed for I3 in the leake current plot (attachment 19) 12:10 No progress on the target so far ... Noticed in aida03 and aida04 several missing channels in the stats spectrum browser (attachment 20) ASIC check load several times on aida03 and aida04. All stats and histograms zero'd 13.37 Target now available - will attempt beam + target New data file G22-88/R16 13.39 per FEE64 Stat spectra 13.43 All histograms and stats zero'd multiple channels missing aida03 and aida04 ( and possibly other FEE64s too )
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown